28.nCr = 30 Normal distribution was discovered by 31 nPr is equal to
32.Number of heads in tossing 3 coins is 33.Number of insured homes against fire is
34.Poisson distribution has/have parameter
35.Probability function can never be 36.Probability function of binomial distribution is
37.Probability of success + probability of failure is equal to
38.Probability of success in binomial distribution
39.Probability of x can never be 40.Probability of x lies between 41.q + p =
1
Expert's answer
2020-07-06T18:52:35-0400

Solution:

28."nCr= \frac{n!}{r!(n-r)!)}"


  1. normal distribution is discovered by Carl Friedrich Gauss


32.Number if heads in tossing 3 coins is discrete variable .


31: "nPr= \frac{n!}{(n-r)!}"


33..number of insured homes against fire is discrete variable.

34.Poisson distribution has parameter lamda.

35 Negative

36.pdf of binomial distribution = "\binom{n}{r}p^{r}q^{n-r}"

37.1

38.probability of success in binomial distribution is p remains constant for each trial.

39.probability of x can never be less than 0 or more than 1.

40.it lies between 0 and 1.

41. p+q=1